OpenStudy (anonymous):

Carl can paint a room 3 hours faster than Jennifer can. If they work together, they can complete the job in 2 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jennifer to complete this job on her own.

OpenStudy (anonymous):

Let x be how long it takes Jennifer to do job, then 1/x+ (1/(x-3))=1/2, So x=6

OpenStudy (anonymous):